OpenStudy (anonymous):

Find the standard form of the equation of the parabola with a focus at (0, -2) and a directrix at y = 2.

OpenStudy (anonymous):

focus (0,p) =(0,-2), hence directrix p =-2 formula x^2=4p y , replace p =-2, isolate y to get the answer

OpenStudy (anonymous):

y = -8/x^2?

OpenStudy (anonymous):

I don't have that option

OpenStudy (anonymous):

-x^2/8 how about this?

OpenStudy (anonymous):

Yes